# Can a Lightweight Automated AI Pipeline Solve Research-Level Mathematical Problems?

### v1

@@ Abstract
Large language models (LLMs) have recently achieved remarkable success in generating rigorous mathematical proofs, with “AI for Math” emerging as a vibrant field of research [7]. While these models have mastered competition-level benchmarks like the International Mathematical Olympiad [5, 4] and show promise in research applications through auto-formalization [13], their deployment via lightweight, natural-language pipelines for research problems remains underexplored. In this work, we demonstrate that next-generation models (e.g., Gemini 3 Pro, GPT-5.2 Pro), when integrated into a streamlined automated pipeline optimized for citation-based verification, can solve sophisticated research-grade problems. We evaluate our pipeline on two novel datasets: (1) the ICCM [6] problem sets (comparable to the S.-T. Yau College Student Mathematics Contest [9]) proposed by leading mathematicians [10], and (2) the “First Proof” problem set [1], consisting of previously unpublished research questions. Our pipeline generated candidate proofs for all problems in the first two ICCM sets and the “First Proof” set. The solutions for the first two ICCM sets and Problem 4 of the “First Proof” set have been fully verified by our team. All generated proofs have been submitted to the official organization, and our generated results are publicly available at https://github.com/ml1301215/question_sets-test_results. We plan to open-source the complete pipeline methodology in due course.
@@ 4.1 The Verification Bottleneck
The primary challenge has shifted from proof generation to efficient verification. While our pipeline produced candidate solutions for the entire “First Proof” set in minutes, the rigorous verification of a single problem (Problem 4) required hours. This asymmetry underscores an urgent need for more sophisticated AI-assisted verification tools—whether through advances in formal methods, interactive semi-formal interfaces, or explainable reasoning frameworks—to keep pace with the accelerating speed of AI-generated mathematics.

### v2

@@ Abstract
Large language models (LLMs) have recently achieved remarkable success in generating rigorous mathematical proofs, with “AI for Math” emerging as a vibrant field of research [7]. While these models have mastered competition-level benchmarks like the International Mathematical Olympiad [5, 4] and show promise in research applications through auto-formalization [13], their deployment via lightweight, natural-language pipelines for research problems remains underexplored. In this work, we demonstrate that next-generation models (e.g., Gemini 3 Pro, GPT-5.2 Pro), when integrated into a streamlined automated pipeline optimized for citation-based verification, can solve sophisticated research-grade problems. We evaluate our pipeline on two novel datasets: (1) the ICCM [6] problem sets (comparable to the S.-T. Yau College Student Mathematics Contest [9]) proposed by leading mathematicians [10], and (2) the “First Proof” problem set [1], consisting of previously unpublished research questions. Our pipeline generated candidate proofs for all problems in the first two ICCM sets and the “First Proof” set. The solutions for the first two ICCM sets and Problem 4 of the “First Proof” set have been fully verified by our team. All generated proofs have been submitted to the official organization, and our generated results are publicly available at https://github.com/ml1301215/question_sets-test_results. We have open-sourced the code and developed a user-friendly UI for this workflow, accessible at https://github.com/ml1301215/research-math-assistant.
@@ 4.1 The Verification Bottleneck
The primary challenge has shifted from proof generation to efficient verification. While our pipeline produced candidate solutions for the entire “First Proof” set in minutes, the rigorous verification of a single problem (Problem 4) required hours. This asymmetry underscores an urgent need for more sophisticated AI-assisted verification tools—whether through advances in formal methods, interactive semi-formal interfaces, or explainable reasoning frameworks to keep pace with the accelerating speed of AI-generated mathematics.